Q1:Consider a level flight turn of a propeller-driven airplane in which the airplane makes a complete circular turn in 1.8 min. If the plane's speed is 200 m/s, what is the radius of the circle?
What is the centripetal acceleration of the plane?
1
Expert's answer
2012-03-30T08:28:33-0400
Let's find the length of a circle:

L = V·T = 200*108 = 21600 m.

We know that the lenght of the circle can be represanted as L = 2πR.
So, the radius of this circilar is

R = L/(2π) = 21600/(2*3.14) ≈ 3439.4904 m.

The centripetal acceleration of the plane is

Ac = V²/R = 200²/3439.4904 ≈ 11.630 m/s².